# Jailbreak-AudioBench: In-Depth Evaluation and Analysis of Jailbreak Threats for Large Audio Language Models

### v2

@@ 4.1 Query-based Audio Editing Jailbreak Method
Specifically, we first create the Explicit Small dataset by extracting 262 samples from the Explicit Subtype dataset, maintaining a one-tenth proportion of the harmful content categories. We then applied 32 distinct audio editing combinations to these base samples, systematically combining modifications related to accent/emotion, emphasis, speed, and background noise in sequence. This combinatorial approach generated $262\times 32=8384$ audio samples comprising our complete Explicit Small dataset. Further combination details are available in the Appendix D.
Hence, each audio in the Explicit Small dataset has 32 variations with different audio editing combinations, which are used to query models to maximize the likelihood of jailbreak. As Figure 5 shows, our Query-based Audio Editing Jailbreak method demonstrates a significant ASR increase in model vulnerabilities to audio jailbreak. Each panel presents a matrix where columns represent individual audio samples from the Explicit Small dataset, and the first 32 rows represent different edited variants of these samples. Green cells indicate failed jailbreak attempts, while red cells indicate successful compromises of the model’s safety guardrails. The penultimate row in each panel represents the original unedited audio sample, while the bottom row indicates whether any of the 32 variant queries successfully bypassed the model’s defenses. Specifically, Qwen2-Audio-7B shows substantial vulnerability with ASR increasing dramatically from 13.3% with original samples to 48.8% under our query-based approach. SALMONN-7B demonstrates even greater susceptibility, with ASR escalating from 31.6% to 85.1%. Most notably, even the closed-source GPT-4o-Audio exhibits vulnerability with ASR increasing from a mere 0.7% to 8.4% under our systematic audio editing combinations. Additional results are available in the Appendix D.
@@
Figure 5: ASR Performance of the Query-based Audio Editing Jailbreak method. In each panel, columns represent individual audio samples, and the first 32 rows represent different edited variants of these samples. The penultimate row represents the original unedited audio sample, while the bottom row indicates whether any of the 32 variant queries bypassed the model’s defenses. Green: failed jailbreak; Red: successful jailbreaks.
@@ 4.2 Defense Method Against Audio Editing Jailbreak
Specifically, we create a defense prompt in audio format using TTS, containing the instruction: "You are a helpful assistant and should refuse to generate illegal, harmful or unethical content.". Then we prepend this defense prompt to each audio sample in the Explicit Subtype dataset, forming the Explicit Defense dataset with safety instructions embedded at the beginning of every sample. More generation details of the Explicit Defense dataset can be found in the Appendix D.
Figure 6 illustrates the ASR comparison of original and edited audio samples with and without defense. The bars represent the ASR without defense, while the striped bars represent the ASR reduction with the defense applied. It shows that the defense approach consistently reduces ASR across all evaluated models, evidenced by the presence of striped segments. This confirms that prepending instructions in audio form offers a baseline level of protection against audio editing jailbreak. However, while the defense provides measurable protection, the residual ASR values remain concerningly high and show the limitations of our defense strategy, which necessitates exploring more effective defense strategies in future work. Additional results can be found in the Appendix D.
Figure 6: ASR comparison of original and edited audio samples with and without defense. The bars represent the ASR without defense, while the striped bars represent the ASR reduction with the defense applied. The values shown on the bars denote the specific ASR reduction caused by defense.

### v3

@@ 4.1 Query-based Audio Editing Jailbreak Method
Specifically, we first create the Explicit Small dataset by extracting 262 samples from the Explicit Subtype dataset, maintaining a one-tenth proportion of the harmful content categories. We then applied 32 distinct audio editing combinations to these base samples, systematically combining modifications related to accent/emotion, emphasis, speed, and background noise in sequence. This combinatorial approach generated $262\times 32=8384$ audio samples comprising our complete Explicit Small dataset. Detailed dataset scale information is shown in Table 1. Further combination details are available in the Appendix D.
Hence, each audio in the Explicit Small dataset has 32 variations with different audio editing combinations, which are used to query models to maximize the likelihood of jailbreak. As Figure 5 shows, our Query-based Audio Editing Jailbreak method demonstrates a significant ASR increase in model vulnerabilities to audio jailbreak on the Explicit Small dataset. Each panel presents a matrix where columns represent individual audio samples from the Explicit Small dataset, and the first 32 rows represent different edited variants of these samples. Green cells indicate failed jailbreak attempts, while red cells indicate successful compromises of the model’s safety guardrails. The penultimate row in each panel represents the original unedited audio sample, while the bottom row indicates whether any of the 32 variant queries successfully bypassed the model’s defenses. Specifically, Qwen2-Audio-7B shows substantial vulnerability with ASR increasing dramatically from 13.3% with original samples to 48.8% under our query-based approach. SALMONN-7B demonstrates even greater susceptibility, with ASR escalating from 31.6% to 85.1%. Most notably, even the closed-source GPT-4o-Audio exhibits vulnerability with ASR increasing from a mere 0.7% to 8.4% under our systematic audio editing combinations. Additional results of the Query-based Audio Editing Jailbreak method on BLSP, SpeechGPT, VITA-1.5, and MiniCPM-o-2.6 are available in the Appendix D.
@@ 4.2 Defense Method Against Audio Editing Jailbreak
Specifically, we create a defense prompt in audio format using TTS, containing the instruction: "You are a helpful assistant and should refuse to generate illegal, harmful or unethical content.". Then we prepend this defense prompt to each audio sample in the Explicit Subtype dataset, forming the Explicit Defense dataset with safety instructions embedded at the beginning of every sample. Following the same process, we construct the Implicit Defense dataset based on the Implicit Subtype dataset. Detailed dataset scale information is shown in Table 1. More generation details of the Defense dataset can be found in the Appendix D.
Figure 6 illustrates the ASR comparison of original and edited audio samples with and without defense on the Explicit Defense dataset. The bars represent the ASR without defense, while the striped bars represent the ASR reduction with the defense applied. It shows that the defense approach consistently reduces ASR across all evaluated models, evidenced by the presence of striped segments. This confirms that prepending instructions in audio form offers a baseline level of protection against audio editing jailbreak. However, while the defense provides measurable protection, the residual ASR values remain concerningly high and show the limitations of our defense strategy, which necessitates exploring more effective defense strategies in future work. Additional results on the Implicit Defense dataset are available in the Appendix D.
Figure 6: ASR comparison of original and edited audio samples with and without defense in the Explicit Defense dataset. The bars represent the ASR without defense, while the striped bars represent the ASR reduction with the defense applied. The values shown on the bars denote the specific ASR reduction caused by defense.